# ADMM Based Semi-Structured Pattern Pruning Framework For Transformer

### v2

@@ Abstract.
Abstract:NLP(natural language processsing) has achieved great success through the transformer model.However, the model has hundreds of millions or billions parameters,which is huge burden for its deployment on personal computer or small scale of server.To deal with it, we either make the model’s weight matrix relatively sparser, or compress attention layer. Pattern pruning ,one of the most important pruning methods, permits selecting fixed number of parameters in each divided pattern block and prunes it. However, the effect of pattern pruning is strictly limited by the sparsity within a region of weights in each layer. In this paper,we first introduced Alternating Direction Method of Multipliers(ADMM) based pattern pruning framework to reshape the distribution of activation map. Specifically, we propose to formulate the pattern pruning on transformer as a constrained optimization and use ADMM to optimize the problem. In this way, the initial dense feature maps is transformed to rather regionally sparsified ones.Therefore, we can then achieve higher compression ratio with better performance based on pattern pruning method. Additionally, this paper provides a theoretical derivations of the ADMM with local sparsity. Finally, we also extend the proposed ADMM based framework on quantization to demonstrate its generalization and use SR-STE to avoid gradient vanishing problem. We conduct extensive experiments on classification tasks over GLUE datasets. Significantly, we achieve 50% percent compression ratio while maintaining overall score 80.1 on GLUE dataset.
model compression, pruning, software, quantization, ADMM
@@ 1. Introduction
Pruning is one of the great solutions of model sparsification, which can be categorized to unstructured pruning(cNN-pruning-with-reduction, ; iterative-pruning, ; Neural-net-pruning-why-and-how, ),, structured pruning(strucurepruning, ; structured-2, ; struturepruning-llm, ) and semi-structured pruning(NxMTransformer, ; dnn_patternpruning, ; n:M-sparsity, ; PP-Transformer, ). The basic theory of pruning is remove the parameter which is around zero meaning low importance. Unstructured pruning directly removes these individual parameters, which usually maintains the best accuracy but is unfriendly to the hardware implementation because of its irregularity. Structured pruning(strucurepruning, ) remove parameters in groups(e.g, two rows or columns), which significantly ease the hardware burden but may reduce the accuracy. Semi-structured pruning maintains the accuracy and the sparsity of specified granularity, reaching a good balance in terms of flexibility and efficiency.
@@
We gain the motivation to choose ADMM to optimize pattern pruning is for the weight matrix of Bert(and other transformer models) are with poor pattern sparsity, which refers to the submatrixs divided by pattern pruning are unbalanced for distribution on the aspect of sparsity. Thus, the effect of ADMM can alleviate the phenomenon and change the polarized distribution in this pre-pruning step into the shape that frankly spread around 50% percent of sparsity while maintaining the accuracy, as the Fig.2. shows. That’s the core of ADMM application to boost the efficiency of pruning. Meanwhile, we can gain better effect and accuracy because ADMM can also be applied to quantization jointly.
In this paper, we present a systematic ADMM(admmref, ) based pattern pruning Transformer framework with the extension of weight quantization. We make the following contributions:(1)To our best knowledge, it’s the first time applying the pattern pruning on both FFN and Attention layer to achieve the bigger compression ratio. We first introduce Sparse-refined straight-through estimator (SR-STE(SR-STE, )) in Transformer pattern pruning training process to deal with the vanishing gradient and accuracy degradation problem.(2)We first respectively formulate the pattern pruning problem and model quantization as a constrained nonconvex problem. We adopt the Alternating Direction Method of Multipliers(ADMM)(admmref, ) to systematically solve pattern pruning optimization problem on Transformer. By using ADMM, the optimization problem is then decomposed into two smaller sub problems which can be solved iteratively. In the pattern pruning problem and quantization problem, one of their sub problems can be solved by stochastic gradient descent and the other can be solved through closed-form solution derived from the equation. (3)With extensive experiments on GLUE dataset, we make detailed analysis of the effect of our framework powered by ADMM and SR-STE are efficient to optimize pattern pruning and basic quantization on whole Transformer model. These optimization methods are complimentary to other compression method including knowledge distillation(knowledgedistilling, ).
@@ 4. Evaluation
Implementation. We conduct experiments on a NVIDIA 3090 GPU with batch size 24, and use Adamw optimizer(ADAMW, ) with the initial learning rate of 7e-5 meanwhile using learning rate scheduler to make adaptive adjustments during the training process. Both pre-training and pattern pruning were trained for 5 epochs. The experimental section uses Bert as the base model for pruning. In order to test the performance on different downstream tasks, we use Bert-base as an example. The data processing and other experimental settings remain consistent with the default settings of Bert. The experimental environment consists of Python 3.9.3, PyTorch 1.13.1, and CUDA 12.3.
@@ 4.1. Main results
We search for the optimal learning rates (7e-5) and weight decay (0.001) to achieve the best possible results. The parameter rho in the ADMM loss is adjusted in the experiment process and finally through the change process of rho, we fix rho as 0.01 and gain the good score on different tasks comprehensively. Table 1 shows the evaluation results for BERT, and the following key observations can be made. First, the basic pruning-based method sparsifies the weights of Transformer blocks but cannot explicitly satisfy the underlying hardware constraints, e.g., the pattern sparsity. While the highest accuracy on downstream tasks was preserved (81.3 vs. 81.8 on average), the obtained sparse weights had a random structure of non-zero weights, which was inefficient and unacceptable. Second, the original pattern pruning method, which was not optimized using the ADMM technique, are demonstrated to have a relatively poor performance, both in individual downstream tasks and on average(78.1%). In contrast, the pattern pruning framework optimized through ADMM demonstrated accuracy levels in various tasks that approached those of the N×M Transformer, which had been optimized by multiple methods. Notably, it achieved scores of 55.4 and 68.8 on the COLA and RTE tasks, surpassing the N×M Transformer. The overall average score reached 80.1, which was significantly higher than that of the pre-optimization pattern pruning method by near 2 percent.
@@ 4.2. Effect on weight distribution
Furthermore, we investigate the impact of our ADMM-based framework on the weight distribution of each layer. Fig.4. illustrates the alteration in the weight distribution of 0th layer of Bert following the training process within our ADMM-based framework. Our framework can effectively tailor to different shape’s weight and according to the division to purposefully make sparsity of each part approach our given sparsity.Here we set it as 50%. In Fig.4., ADMM-optimized weight matrix (the query weight of layer 0) exhibits a notable increase in zero values compared to the original matrix. This demonstrates the effectiveness of the ADMM optimization in achieving a 50% sparsity distribution for pattern pruning.
@@ 4.3. Analysis of results
Our best accuracy of Bert model after optimization, before pruning, achieves 89.7%. After fully pruning the model, the average score of Bert model can achieve 80.1% while maintaining 50% percent of sparsity. On specific task, COLA and RTE, it can outperform the N×M Transformer, which is an excellent framework nowadays.
Validation and hardware-friendness improvement Our evaluation scores and hardware efficiency are clearly demonstrated in the data and charts. The significant improvements can be attributed to our ADMM optimization framework. Within this framework, we integrated the SR-STE method to prevent gradient vanishing, enhancing the training effect.During pattern pruning, varying sparsity levels in different weight blocks led to substantial discrepancies between the pruned matrices and the original large matrices, either by excessively reducing large values or retaining too many small values in certain blocks, severely affecting validation accuracy. Our ADMM-based framework iteratively optimizes the weight distribution, ensuring that the weights are more evenly distributed to approximately 50% sparsity before pattern pruning. This approach significantly boosts the post-pruning scores, achieving an overall score of 80.1, which is 2 percentage points higher than pre-optimization and only 0.4 percentage points lower than one of the best current sparse frameworks, the N×M Transformer.Additionally, the structural uniformity achieved through our pruning method provides hardware efficiency comparable to the N×M Transformer. By leveraging and improving the existing pattern pruning FPGA hardware framework, we can achieve excellent deployment results.
@@ 5. Future works
According to PP-Transformer(PP-Transformer, ), the performance of pattern pruning(bigger pattern) will drop quickly when sparsity becomes above 80%percent. In our experiment process, we found that after ADMM improvement, the performance with high sparsity, bigger pattern choice pattern pruning significantly can partly deal with that problem, but still with unacceptable performance drop. Therefore, we want to find other optimization method and combine it with ADMM to completely stop it from dropping. On the other hand, we hope we can develop a complete set of hardware framework tailoring to ADMM optimization, which will greatly speed up the deploy and run time on FPGA or ASIC chips.
@@ 6. Conclusions
ADMM can bring the transformer weight distribution to the shape we desire and greatly improve the pruning performance. This paper demonstrate the effectiveness of ADMM on pattern pruning from performance and weight distribution aspects. Furthermore, ADMM based pattern pruning Framework is orthogonal to quantization and other optimization techniques. This framework we proposed can be wildly applied to almost all the models based on the transformer. including Bert,Vit and others.While the framework do introduce a loss and accuarcy gap because of compression, we believe the ADMM based pattern pruning framework will work and keep its importance.

### v3

@@ Abstract.
Abstract:NLP(natural language processsing) has achieved great success through the transformer model.However, the model has hundreds of millions or billions parameters,which is huge burden for its deployment on personal computer or small scale of server.To deal with it, we either make the model’s weight matrix relatively sparser, or compress attention layer. Pattern pruning ,one of the most important pruning methods, permits selecting fixed number of parameters in each divided pattern block and prunes it. However, the effect of pattern pruning is strictly limited by the sparsity within a region of weights in each layer. In this paper,we first introduced Alternating Direction Method of Multipliers(ADMM) based pattern pruning framework to reshape the distribution of activation map. Specifically, we propose to formulate the pattern pruning on transformer as a constrained optimization and use ADMM to optimize the problem. In this way, the initial dense feature maps is transformed to rather regionally sparsified ones.Therefore, we can then achieve higher compression ratio with better performance based on pattern pruning method. Additionally, this paper provides a theoretical derivations of the ADMM with local sparsity. Finally, we also extend the proposed ADMM based framework with SR-STE to demonstrate its generalization and to avoid gradient vanishing problem. We conduct extensive experiments on classification tasks over GLUE datasets. Significantly, we achieve 50% percent compression ratio while maintaining overall score 80.1 on GLUE dataset.
model compression, pruning, software, ADMM
@@ 1. Introduction
Pruning is one of the great solutions of model sparsification, which can be categorized to unstructured pruning(cNN-pruning-with-reduction, ; iterative-pruning, ; Neural-net-pruning-why-and-how, ), structured pruning(strucurepruning, ; structured-2, ; struturepruning-llm, ) and semi-structured pruning(NxMTransformer, ; dnn_patternpruning, ; n:M-sparsity, ; PP-Transformer, ). The basic theory of pruning is remove the parameter which is around zero meaning low importance. Unstructured pruning directly removes these individual parameters, which usually maintains the best accuracy but is unfriendly to the hardware implementation because of its irregularity. Structured pruning(strucurepruning, ) remove parameters in groups(e.g, two rows or columns), which significantly ease the hardware burden but may reduce the accuracy. Semi-structured pruning maintains the accuracy and the sparsity of specified granularity, reaching a good balance in terms of flexibility and efficiency.
@@
We gain the motivation to choose ADMM to optimize pattern pruning is for the weight matrix of Bert(and other transformer models) are with poor pattern sparsity, which refers to the submatrixs divided by pattern pruning are unbalanced for distribution on the aspect of sparsity. Thus, the effect of ADMM can alleviate the phenomenon and change the polarized distribution in this pre-pruning step into the shape that frankly spread around 50% percent of sparsity while maintaining the accuracy. That’s the core of ADMM application to boost the efficiency of pruning.
In this paper, we present a systematic ADMM(admmref, ) based pattern pruning Transformer framework. We make the following contributions:(1)To our best knowledge, it’s the first time applying the pattern pruning on both FFN and Attention layer to achieve the bigger compression ratio. We first introduce Sparse-refined straight-through estimator (SR-STE(SR-STE, )) in Transformer pattern pruning training process to deal with the vanishing gradient and accuracy degradation problem.(2)We first respectively formulate the pattern pruning problem as a constrained nonconvex problem. We adopt the Alternating Direction Method of Multipliers(ADMM)(admmref, ) to systematically solve pattern pruning optimization problem on Transformer. By using ADMM, the optimization problem is then decomposed into two smaller sub problems which can be solved iteratively. In the pattern pruning problem, one of their sub problems can be solved by stochastic gradient descent and the other can be solved through closed-form solution derived from the equation. (3)With extensive experiments on GLUE dataset, we make detailed analysis of the effect of our framework powered by ADMM and SR-STE are efficient to optimize pattern pruning on whole Transformer model. These optimization methods are complimentary to other compression method including knowledge distillation(knowledgedistilling, ).
@@ 4. Evaluation
Implementation. We conduct experiments on a NVIDIA 3090 GPU with batch size 24, and use Adamw optimizer(ADAMW, ) with the initial learning rate of 7e-5 meanwhile using learning rate scheduler to make adaptive adjustments during the training process. The weight decay is set to 0.001. Both pre-training and pattern pruning were trained for 5 epochs. The experimental section uses Bert as the base model for pruning. In order to test the performance on different downstream tasks, we use Bert-base as an example. The data processing and other experimental settings remain consistent with the default settings of Bert. The experimental environment consists of Python 3.9.3, PyTorch 1.13.1, and CUDA 12.3.
@@ 4.1. Main results
In the experiment, we iteratively adjusted the value of $\rho$ and ultimately determined that setting $\rho$ to 0.01 yielded the best results across all tasks. Table 1 presents the results of different methods on BERT pruning, and we made the following observations: First, the ADMM(unstructured), a basic pruning-based method, sparsifies the weights of Transformer blocks but cannot explicitly satisfy the underlying hardware constraints, e.g., the pattern sparsity. While the highest accuracy on downstream tasks was preserved (81.3% on average compared to the baseline’s 81.8%), the resulting sparse weights exhibited a random structure of non-zero weights. This random structure was inefficient and considered unacceptable. Second, the pattern pruning method (without ADMM optimization) demonstrated relatively poor performance, both in individual downstream tasks and on average (78.1%). This indicates that the method’s effectiveness was limited without utilizing the ADMM technique. In contrast, our method, pattern pruning optimized with ADMM, achieved superior results across various tasks, even approaching the performance of the N×M Transformer. It is worth noting that the N×M Transformer has undergone extensive optimization. Furthermore, our method is more flexible as it allows for setting the sparsity level for each layer, and even the distribution of each individual weight matrix. Notably, it achieved the best scores of 55.4% and 68.8% on the COLA and RTE tasks, surpassing the N×M Transformer. The overall average score reached 80.1%, which was significantly higher than that of the pre-optimization pattern pruning method by near 2 percents.
@@ 4.2. Effect on weight distribution
Furthermore, we investigate the impact of our ADMM-based framework on the weight distribution of each layer. Fig.4 illustrates the distribution changes of weights in the 0th layer of BERT when before/after using the ADMM-based framework. Our framework can effectively tailor to different shape’s weight and according to the division to purposefully make sparsity of each part approach our given sparsity. Here we set it as 50%. In Fig.4., ADMM-optimized weight matrix (the query weight of layer 0) exhibits a notable increase in zero values compared to the original matrix. This demonstrates the effectiveness of the ADMM optimization in achieving a 50% sparsity distribution for pattern pruning.
@@ 4.3. Analysis of results
Our best accuracy of Bert model after optimization, before pruning, achieves 81.8% on average. After fully pruning the model, the average score of Bert model can achieve 80.1% while maintaining 50% percent of sparsity under ADMM optimization. On specific task, COLA and RTE, it can outperform the N×M Transformer, which is an excellent framework.
Validation and hardware-friendness improvement Our evaluation scores and hardware efficiency are clearly demonstrated by quantitative experiments. The significant improvements can be attributed to our ADMM optimization framework. Within this framework, we integrated the SR-STE method to prevent gradient vanishing, enhancing the training effect. During pattern pruning, varying sparsity levels in different weight blocks led to substantial discrepancies between the pruned matrices and the original large matrices, severely affecting validation accuracy. Our ADMM-based framework iteratively optimizes the weight distribution, ensuring that the weights are more evenly distributed to approximately 50% sparsity before pattern pruning. This approach significantly boosts the post-pruning scores, achieving an overall score of 80.1, which is 2 percentage points higher than pre-optimization and only 0.4 percentage points lower than one of the best current sparse frameworks, the N×M Transformer.Additionally, the structural uniformity achieved through our pruning method provides hardware efficiency comparable to the N×M Transformer. By leveraging and improving the existing pattern pruning FPGA hardware framework, we can achieve excellent deployment results.
@@ 5. Future works
According to PP-Transformer(PP-Transformer, ), the performance of pattern pruning with the bigger division (e.g. 8x8) will drop quickly when sparsity becomes above 80%percent. In our experiment, we observed that even after applying ADMM improvement, the performance degradation with high sparsity and larger pattern choice pattern pruning remained minimal but still exhibited an unacceptable drop in performance. Therefore, we want to find other optimization method and combine it with ADMM to completely stop it from dropping. On the other hand, we hope we can develop a complete set of hardware framework tailoring to ADMM optimization, which will greatly speed up the deploy and run time on FPGA or ASIC chips.
@@ 6. Conclusions
ADMM can reshape the weight distribution of the transformer to the desired form and significantly enhance the pruning performance. This paper demonstrate the effectiveness of ADMM on pattern pruning from performance and weight distribution aspects. Furthermore, ADMM based pattern pruning framework is orthogonal to many other optimization techniques. This framework we proposed can be widely applied to almost all the models based on the transformer, including Bert, ViT and others. While the framework introduces an accuracy gap due to compression, we firmly believe that the ADMM-based pattern pruning framework will continue to be effective and maintain its significance.